OpenStudy (anonymous):

Prove that for n>=1, n choose r = n choose (n-r)

OpenStudy (zarkon):

\[{n \choose r}=\frac{n!}{r!(n-r)!}=\frac{n!}{(n-r)!r!}=\frac{n!}{(n-r)!(n-(n-r))!}={n\choose n-r}\]

OpenStudy (anonymous):

thanks, this whole time I had the equation wrong. lol